# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

### v4

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a cutting-edge technique for real-time radiance field rendering, offering state-of-the-art performance in terms of both quality and speed. 3DGS models a scene as a collection of three-dimensional Gaussians, or "splats," with additional attributes optimized to conform to the scene’s geometric and visual properties. Despite its advantages in rendering speed and image fidelity, 3DGS is limited by its significant storage and memory demands. These high demands make 3DGS impractical for mobile devices or headsets, reducing its applicability in important areas of computer graphics. To address these challenges and advance the practicality of 3DGS, this state-of-the-art report (STAR) provides a comprehensive and detailed examination of compression and compaction techniques developed to make 3DGS more efficient. We categorize current approaches into compression techniques, which aim at achieving the highest quality at minimal data size, and compaction techniques, which aim for optimal quality with the fewest Gaussians. We introduce the basic mathematical concepts underlying the analyzed methods, as well as key implementation details and design choices. Our report thoroughly discusses similarities and differences among the methods, as well as their respective advantages and disadvantages. We establish a consistent standard for comparing these methods based on key performance metrics and datasets. Specifically, since these methods have been developed in parallel and over a short period of time, currently, no comprehensive comparison exists. This survey, for the first time, presents a unified standard to evaluate 3DGS compression techniques. To facilitate the continuous monitoring of emerging methodologies, we maintain a dedicated website that will be regularly updated with new techniques and revisions of existing findings https://w-m.github.io/3dgs-compression-survey/. Overall, this STAR provides an intuitive starting point for researchers interested in exploring this rapidly growing field. By comprehensively categorizing and evaluating existing compression and compaction strategies, our work advances the understanding and practical application of 3DGS in computationally constrained environments.
@@ 1 Introduction
3D Gaussian Splatting (3DGS)[18] has emerged as an explicit alternative to NeRFs, which delivers quality close to the state-of-the-art NeRF approach of Zip-NeRF[5], while allowing explicit control over scene elements. 3DGS uses Gaussian Splatting to represent a scene by distributing a collection of Gaussian ellipsoids (or splats) throughout the volume. These splats approximate surfaces and volumes, providing an efficient means of rendering and allowing for more straightforward scene manipulation and editing. 3DGS builds on earlier computer graphics work on splatting techniques [44], such as EWA splatting[52], which laid the foundation for point-based rendering methods. A key advantage of 3DGS is its ability to handle complex visual effects, such as view-dependent lighting and transparency, which are vital for achieving accurate and visually convincing representations of real-world scenes. As an explicit representation involving Gaussian ellipsoids, 3DGS is editable and can be composed with other representations, such as textured meshes, which are widely used across industries like gaming and film due to their flexibility, editability, and compatibility with existing rendering engines. By supporting features like transparency and view-dependent effects, 3DGS enhances the versatility of 3D scene recording.
@@
Moreover, 3DGS supports real-time rendering on high-end GPUs, creating an optimistic outlook for broader availability of applications on mobile and VR devices in the near future. Its simplicity also contributes to its accessibility: the 3DGS format uses a simple renderer with multiple open implementations and a basic .ply file format[42], which can be read and written by libraries or implementations in many programming languages, making 3DGS accessible from the day they were published. This simplicity has driven rapid adoption across various fields[46], making 3DGS an attractive option for both researchers and practitioners. However, early implementations of 3DGS had significant file size challenges [39]. Storing all attributes with full precision often resulted in files of several gigabytes, which drove a wave of research focused on compression techniques. Compression of 3DGS files provides several benefits: it allows for faster transmission over slower connections, enables rendering on lower-end devices by reducing memory requirements, and supports the creation of larger, more complex scenes, particularly in video game development. There is thus a strong motivation for compression, leading to an ongoing race to discover the most effective techniques that balance realism and computational efficiency.
@@
The goal of compact scene representation is to achieve an optimal balance between visual realism and efficient storage. While NeRFs perform compression as an integral part of their neural network-based representation, 3DGS requires explicit densification - adding more Gaussian splats to fill in details - and pruning - removing redundant splats to optimize memory usage and performance. Despite these challenges, the explicit nature of 3DGS makes it particularly effective for recording real-world scenes, especially with its ability to capture subtle lighting effects and transparency.
@@
This survey provides an overview of the different techniques developed to achieve a compact 3DGS representation, summarizing numerous parallel efforts and evaluating their effectiveness. By examining these methods, the survey guides future research and applications, helping the community identify successful approaches and areas that require further exploration. Many of the techniques used for compressing 3DGS are adaptations of classical methods, tailored for Gaussian splats, while others introduce entirely new approaches to reduce data size.
The community has been closing the gap in compression efficiency between 3DGS and NeRFs, which may currently represent an optimal benchmark. By presenting a comprehensive overview, we assist the research community in merging successful compression techniques into an optimal solution for 3DGS. Ultimately, developing a simple, straightforward standard for compressed 3DGS will foster broader adoption in the computer graphics community, making it more versatile for different applications. Such a standard will ensure compatibility, ease of use, and encourage further innovation, helping 3DGS become as ubiquitous and practical as textured meshes across a range of use cases.
@@
In this state-of-the-art report, we focus on optimization techniques for 3D Gaussian Splatting (3DGS) representations, aiming to optimize memory usage while preserving visual quality and real-time rendering speed. We focus on compression and compaction methods. We provide a comprehensive comparison of various compression techniques, with quantitative results for the most commonly used datasets summarized in a tabulated format. Specifically, We aim to ensure transparency and create a basis for reproducibility of the included approaches. Additionally, we offer a brief explanation of each pipeline and compare and discuss the main compression and compaction approaches. Rather than covering all existing 3DGS methods and applications, our focus is specifically on techniques to optimize 3DGS representations for size or memory footprint; for a broader overview of 3DGS methods and applications, we refer readers to [46, 13]. While we include common approaches shared between neural radiance field (NeRF)[33] compression and 3DGS compression, we direct readers to [24, 9] for NeRF-specific compression methods.
@@ 2 Fundamentals of 3D Gaussian Splatting (3DGS)
3D Gaussian Splatting[18] introduces a novel approach to real-time radiance field rendering, achieving state-of-the-art performance quality and rendering speed. The technique involves depicting a scene as an ensemble of 3D Gaussians, referred to as “splats”, which are optimized to align with the scene’s geometry and visual characteristics.
The training of a static 3D Gaussian Splatting scene utilizes a collection of images as input, in conjunction with calibrated cameras derived from Structure from Motion (SfM), which yield a sparse initial point cloud. For each point within the resulting sparse cloud, a 3D Gaussian distribution is initialized as follows:
$G(\mathbf{x})=\alpha\cdot e^{-\frac{1}{2}(\mathbf{x-\mu})^{T}\Sigma^{-1}(\mathbf{x-\mu})}$ (1)
Each 3D Gaussian is characterized by its 59 attributes:
@@
  • Covariance matrix ($\Sigma$): The covariance matrix, representing orientation and size of the splat, can be factorized into the scaling $S$ along the x, y, and z axes (3 attributes) and rotation $R$, a quaternion, which has 4 components. An exponential activation is used on the scaling parameters. (7 attributes)
@@
  • Opacity ($\alpha$): A single scalar value. A sigmoid activation is used to constrain the values to range $[0,1)$. (1 attribute)
@@
  • Color ($c$): View-dependent color is represented by Spherical Harmonics (SH) coefficients. The first degree of spherical harmonics is required to give splats a color, split in 3 channels (R, G, B). These are denoted $f\_dc\_\{0,1,2\}$ in the .ply point cloud format. Additional view-dependent colors are supported through higher degrees of SH. 3DGS proposes to use degree 3, thus 15 additional coefficients per color channel. (3 rgb x 16 coefficients = 48 attributes)
Figure 1: Histograms of the opacity and the first scaling attributes for all splats of the bicycle scene (as trained by 3DGS[18]).
Figure 1 shows histograms of the opacity and the first scaling attributes for all splats of the bicycle scene (as trained by 3DGS[18]). Before passing the values to the renderer, their activation functions are applied: sigmoid for the opacity, and exponential for the scale. We can see a peak of very low opacity splats, demonstrating potential candidate splats to be removed from the scene (see Figure 5). In the scaling histogram, we can see that there are very few large splats, an effect from applying the adaptive density control. Additional histograms for the same scene are shown in the Appendix A.
@@
The rendering of 3D Gaussian splats is achieved by their projection from a three-dimensional space onto a two-dimensional image plane. Each three-dimensional Gaussian is transformed into a two-dimensional Gaussian whose footprint is derived from its covariance matrix and the camera’s view transformation parameters. For each pixel, colors are aggregated via alpha-blending, with contributions from each Gaussian blended according to their depth order. Ultimately, the pixel’s color, denoted as C, is determined as demonstrated in Eq. (2):
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (2)
@@
In this context, $N$ denotes the set of depth-sorted Gaussians that overlap at the pixel, $c_{i}$ is the color and $\alpha_{i}$ is the opacity of the i-th Gaussian. This equation ensures that Gaussians are blended in the appropriate sequence, with those situated closer exerting a more significant impact on the resulting pixel color. Throughout the training or optimization phase, the position, size (as represented by the covariance matrix), opacity, and color of the Gaussian functions are systematically refined to optimally correspond to the input views. Differentiable rendering is used to compute gradients, which allows the adjustment of Gaussian parameters so that the rendered images align with the training images.
@@ 3 Fundamentals of Compression and Compaction for 3DGS
While Gaussian Splatting scenes can efficiently be rendered in real-time, they typically have high demands on memory consumption, requiring efficient compression to enable storage of large environments. For other media data like images, video, or audio different coding strategies exist to significantly reduce data size. Usually, coding methods are distinguished between lossless and lossy coding. For lossless coding, redundancy is reduced by exploiting the different probabilities of the individual symbols. Such entropy encoding can be achieved, e.g., with Huffman [16] or arithmetic coding [1]. For correlated sources, coding efficiency can be further increased by appropriate prediction or transformation of symbols prior to entropy coding. However, depending on source statistics, lossless coding is often restricted to moderate compression around a factor of 2. Much higher coding efficiency can be expected when tolerating small deviations in the decoded data. Lossy compression targets at removing irrelevance in the data that cannot perceived by humans but would need additional bits for encoding. Quantization of the predicted or transformed symbols to fewer code words is a standard approach, either for scalar values or an entire vector as in vector quantization (see Sec. 3.1). The goal is then to adjust and shift the quantization error such that is does not become visible.
Figure 2: A correlation heatmap for attributes of all splats of the bicycle scene (as trained by 3DGS[18]).
As an example Figure 2 shows a correlation heatmap for attributes of all splats of the bicycle scene, as provided by 3DGS[18]. The color channels of the base colors of the splats ($f\_dc$) are nearly perfectly correlated, likely due to the dominance of luminance over chrominance changes in this natural scene. This significant correlation suggests the feasibility of jointly encoding the base color channels. Further correlation exists between the individual scale attributes and also between scale and opacity, which can be exploited in compression. A full correlation map including all spherical harmonics attributes can be found in the Appendix A
Beyond standard coding techniques, Gaussian Splatting offers many chances to modify data to support more efficient coding. The reason lies in the different ambiguities within the Gaussian Splatting structure. Different arrangements of splats can lead to the same visual appearance while differing in their compressibility. This can be exploited in considering additional losses during training and scene optimization to limit bit-rate while keeping visual quality. In the following sections, we will first present several concepts for compression and compaction of Gaussian Splatting scenes while Section 4 will dive into details of individual approaches from the current state of the art.
@@ 3.1 Vector Quantization
The quantization process can either be performed on the entire multi-dimensional vectors simultaneously or individual dimensions of the data space can be treated as separate quantization tasks. The latter case allows for more flexible handling of different data attributes, especially for applications where different properties of the data require different levels of precision and have different distributions and levels of redundancy. In essence, vector quantization algorithms improve a small set of vectors to represent a larger set of vectors under some optimization criterion. This results in significant compression gains, making VQ highly efficient when the data exhibits redundancy or for data-rich applications where a slight loss of precision is acceptable in exchange for dramatic reductions in data size. Hence, these approaches play a crucial role in various domains, from signal processing and image compression to the emerging field of 3D data representation to compress complex datasets while maintaining a close approximation to the original. Especially, it is highly effective for compressing 3D Gaussian splatting (3DGS) data, where attributes like positions, radii, colors, or spherical harmonics are often highly redundant and exhibit natural clustering tendencies.

### v5

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a cutting-edge technique for real-time radiance field rendering, offering state-of-the-art performance in terms of both quality and speed. 3DGS models a scene as a collection of three-dimensional Gaussians, with additional attributes optimized to conform to the scene’s geometric and visual properties. Despite its advantages in rendering speed and image fidelity, 3DGS is limited by its significant storage and memory demands. These high demands make 3DGS impractical for mobile devices or headsets, reducing its applicability in important areas of computer graphics. To address these challenges and advance the practicality of 3DGS, this state-of-the-art report (STAR) provides a comprehensive and detailed examination of compression and compaction techniques developed to make 3DGS more efficient. We classify existing methods into two categories: compression, which focuses on reducing file size, and compaction, which aims to minimize the number of Gaussians. Both methods aim to maintain or improve quality, each by minimizing its respective attribute: file size for compression and Gaussian count for compaction. We introduce the basic mathematical concepts underlying the analyzed methods, as well as key implementation details and design choices. Our report thoroughly discusses similarities and differences among the methods, as well as their respective advantages and disadvantages. We establish a consistent framework for comparing the surveyed methods based on key performance metrics and datasets. Specifically, since these methods have been developed in parallel and over a short period of time, currently, no comprehensive comparison exists. This survey, for the first time, presents a unified framework to evaluate 3DGS compression techniques. To facilitate the continuous monitoring of emerging methodologies, we maintain a dedicated website that will be regularly updated with new techniques and revisions of existing findings:
https://w-m.github.io/3dgs-compression-survey/.
Overall, this STAR provides an intuitive starting point for researchers interested in exploring the rapidly growing field of 3DGS compression. By comprehensively categorizing and evaluating existing compression and compaction strategies, our work advances the understanding and practical application of 3DGS in computationally constrained environments.
@@ 1 Introduction
3D Gaussian Splatting (3DGS)[21] has emerged as an explicit alternative to NeRFs, which delivers quality close to the state-of-the-art NeRF approach of Zip-NeRF[5], while allowing explicit control over scene elements. 3DGS uses Gaussian Splatting to represent a scene by distributing a collection of Gaussian ellipsoids throughout the volume. These Gaussians approximate surfaces and volumes, providing an efficient means of rendering and allowing for more straightforward scene manipulation and editing. 3DGS builds on earlier computer graphics work on splatting techniques [51], such as EWA splatting[59], which laid the foundation for point-based rendering methods. A key advantage of 3DGS is its ability to handle complex visual effects, such as view-dependent lighting and transparency, which are vital for achieving accurate and visually convincing representations of real-world scenes. As an explicit representation involving Gaussian ellipsoids, 3DGS is editable and can be composed with other representations, such as textured meshes, which are widely used across industries like gaming and film due to their flexibility, editability, and compatibility with existing rendering engines. By supporting features like transparency and view-dependent effects, 3DGS enhances the versatility of 3D scene recording.
@@
Moreover, 3DGS supports real-time rendering on high-end GPUs, creating an optimistic outlook for broader availability of applications on mobile and VR devices in the near future. Its simplicity also contributes to its accessibility: the 3DGS format uses a simple renderer with multiple open implementations and a basic .ply file format[48], which can be read and written by libraries or implementations in many programming languages, making 3DGS accessible from the day it was published. This simplicity has driven rapid adoption across various fields[52], making 3DGS an attractive option for both researchers and practitioners. However, early implementations of 3DGS had significant file size challenges [44]. Storing all attributes with full precision often resulted in files of several gigabytes, which drove a wave of research focused on compression techniques. Compression of 3DGS files provides several benefits: it allows for faster transmission over slower connections, enables rendering on lower-end devices by reducing memory requirements, and supports the creation of larger, more complex scenes, particularly in video game development. There is thus a strong motivation for compression, leading to an ongoing race to discover the most effective techniques that balance realism and computational efficiency.
@@
The goal of compact scene representation is to achieve an optimal balance between visual realism and efficient storage. While NeRFs perform compression as an integral part of their neural network-based representation, 3DGS requires explicit densification - adding more Gaussian primitives to fill in details - and pruning - removing redundant primitives to optimize memory usage and performance. Despite these challenges, the explicit nature of 3DGS makes it particularly effective for recording real-world scenes, especially with its ability to capture subtle lighting effects and transparency.
@@
This survey provides an overview of the different techniques developed to achieve a compact 3DGS representation, summarizing numerous parallel efforts and evaluating their effectiveness. By examining these methods, the survey guides future research and applications, helping the community identify successful approaches and areas that require further exploration. Many of the techniques used for compressing 3DGS are adaptations of classical methods, tailored for Gaussian splatting, while others introduce entirely new approaches to reduce data size.
The community has been closing the gap in compression efficiency between 3DGS and NeRFs, which may currently serve as an optimal benchmark. By providing a comprehensive overview, we aim to help researchers identify effective strategies and avoid pitfalls, facilitating the development of improved 3DGS compression methods. Establishing a simple and consistent standard for compressed 3DGS will promote broader adoption in the computer graphics community, enhancing its versatility for diverse applications. Such a standard will streamline compatibility and usability, paving the way for 3DGS to achieve the ubiquity and practicality of textured meshes across various use cases.
@@
In this state-of-the-art report, we focus on optimization techniques for 3D Gaussian Splatting (3DGS) representations, aiming to optimize memory usage while preserving visual quality and real-time rendering speed. We focus on compression and compaction methods. We provide a comprehensive comparison of various compression techniques, with quantitative results for the most commonly used datasets summarized in a tabulated format. Specifically, we aim to ensure transparency and create a basis for reproducibility of the included approaches. Additionally, we offer a brief explanation of each pipeline and compare and discuss the main compression and compaction approaches. Rather than covering all existing 3DGS methods and applications, our focus is specifically on techniques to optimize 3DGS representations for size or memory footprint; for a broader overview of 3DGS methods and applications, we refer readers to [52, 14]. While we include common approaches shared between neural radiance field (NeRF)[38] compression and 3DGS compression, we direct readers to [28, 8] for NeRF-specific compression methods. Furthermore, we acknowledge the inclusion of both peer-reviewed publications from conferences and journals, as well as preprints on arXiv, within our state-of-the-art report, recognizing the rapidly evolving nature of the field of 3D Gaussian splatting.
@@ 2 Fundamentals of 3D Gaussian Splatting (3DGS)
3D Gaussian Splatting[21] introduces a novel approach to real-time radiance field rendering, achieving state-of-the-art performance quality and rendering speed. The technique involves depicting a scene as an ensemble of 3D Gaussian primitives, which are optimized to align with the scene’s geometry and visual characteristics. Each 3D Gaussian is characterized by its 59 attributes:
@@
  • Covariance matrix ($\Sigma$): The covariance matrix, representing orientation and size of the Gaussian primitive, can be factorized into the scaling $S$ along the x, y, and z axes (3 attributes) and rotation $R$, a quaternion, which has 4 components. An exponential activation is used on the scaling parameters. (7 attributes)
@@
  • Opacity ($o$): A single scalar value. A sigmoid activation is used to constrain the values to range $(0,1)$. (1 attribute)
@@
  • Color ($c$): View-dependent color is represented by Spherical Harmonics (SH) coefficients. The first degree of spherical harmonics is required to give splats a diffuse color, split into 3 channels (R, G, B). These are denoted $f\_dc\_\{0,1,2\}$ in the .ply point cloud format. Additional view-dependent colors are supported through higher degrees of SH. 3DGS proposes to use degree 3, thus 15 additional coefficients per color channel. (3 rgb x 16 coefficients = 48 attributes)
@@
The training of a static 3D Gaussian Splatting scene utilizes a collection of images as input, in conjunction with calibrated cameras derived from Structure from Motion (SfM), which yield a sparse initial point cloud. For each point within the resulting sparse cloud, a 3D Gaussian distribution ($G(x)$) is initialized. The rendering of 3D Gaussian primitives, also called splatting, is achieved by their projection from a three-dimensional space onto a two-dimensional image plane. Each three-dimensional Gaussian is transformed into a two-dimensional Gaussian (a splat) whose footprint is derived from its covariance matrix and the camera’s view transformation parameters. For each pixel, colors are aggregated via alpha-blending, with contributions from each splat blended according to their depth order. Ultimately, the pixel’s color, denoted as I, is determined as follows:
$I(x)=\sum_{i\in N}\alpha_{i}(x)c_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}(x)),$ (1)
$\hskip 20.00003pt\alpha(x)=oG(x),\hskip 20.00003ptG(x)=e^{-\frac{1}{2}(\mathbf{x-\mu})^{T}\Sigma^{\prime-1}(x-\mu)}$
@@
In this context, $N$ denotes the set of depth-sorted splats that overlap at a pixel, $c_{i}$ is the color and $\alpha_{i}$ is the contribution of a primitive, that is the product of opacity ($o$) and Gaussian falloff. This ensures that splats are blended in the appropriate sequence, with those situated closer exerting a more significant impact on the resulting pixel color. Throughout the training or optimization phase, the position, size (as represented by the covariance matrix), opacity, and color of the Gaussian functions are systematically refined to optimally correspond to the input views. Differentiable rendering is used to compute gradients, which allows the adjustment of Gaussian parameters so that the rendered images align with the training images.
Figure 1: Histograms of the opacity and the first scaling attributes for all 3D Gaussians of the bicycle scene (as trained by 3DGS[21]).
Figure 1 shows histograms of the opacity and the first scaling attributes for all 3D Gaussians of the bicycle scene (as trained by 3DGS[21]). Before passing the values to the renderer, their activation functions are applied: sigmoid for the opacity, and exponential for the scale. We can see a peak of very low opacity Gaussians, demonstrating potential candidates to be removed from the scene (see Figure 10). In the scaling histogram, we can see that there are very few large Gaussians, an effect from applying the adaptive density control. Additional histograms for the same scene are shown in the Appendix A.
@@ 3 Fundamentals of Compression and Compaction for 3DGS
While Gaussian Splatting scenes can efficiently be rendered in real-time, they come with significant file size demands, necessitating efficient compression for managing large scenes. For other media data like images, video, or audio different coding strategies exist to significantly reduce data size. Usually, coding methods are distinguished between lossless and lossy coding. For lossless coding, redundancy is reduced by exploiting the different probabilities of the individual symbols. Such entropy encoding can be achieved, e.g., with Huffman [18] or arithmetic coding [1]. For correlated sources, coding efficiency can be further increased by appropriate prediction or transformation of symbols prior to entropy coding. However, depending on source statistics, lossless coding is often restricted to moderate compression around a factor of 2. Much higher coding efficiency can be expected when tolerating small deviations in the decoded data. Lossy compression targets at removing irrelevance in the data that cannot perceived by humans but would need additional bits for encoding. Quantization of the predicted or transformed symbols to fewer code words is a standard approach, either for scalar values or an entire vector as in vector quantization (see Sec. 3.1). The goal is then to adjust and shift the quantization error such that is does not become visible.
Figure 2: A correlation heatmap for attributes of all 3D Gaussians of the bicycle scene (as trained by 3DGS[21]).
As an example, Figure 2 shows a correlation heatmap for attributes of all 3D Gaussians of the bicycle scene, as provided by 3DGS[21]. The color channels of the base colors of the splats ($f\_dc$) are nearly perfectly correlated, likely due to the dominance of luminance over chrominance changes in this natural scene. This significant correlation suggests the feasibility of jointly encoding the base color channels. Further correlation exists between the individual scale attributes and also between scale and opacity, which can be exploited in compression. A full correlation map including all spherical harmonics attributes can be found in the Appendix A
Beyond standard coding techniques, Gaussian Splatting provides numerous opportunities to modify data to support more efficient coding. The reason lies in the different ambiguities within the Gaussian Splatting structure. Different arrangements of splats can lead to the same visual appearance while differing in their compressibility. This can be exploited in considering additional losses during training and scene optimization to limit bit-rate while keeping visual quality. Reparameterizing the 3DGS representation offers another avenue for achieving compression without compromising visual quality. In the following sections, we will first present several concepts for compression and compaction of Gaussian Splatting scenes, while Section 4 will dive into details of individual approaches from the current state of the art.
@@ 3.1 Vector Quantization
The quantization process can either be performed on the entire multi-dimensional vectors simultaneously or individual dimensions of the data space can be treated as separate quantization tasks. The latter case allows for more flexible handling of different data attributes, especially for applications where different properties of the data require different levels of precision and have different distributions and levels of redundancy. In essence, vector quantization algorithms improve a small set of vectors to represent a larger set of vectors under some optimization criterion. This results in significant compression gains, making VQ highly efficient when the data exhibits redundancy or for data-rich applications where a slight loss of precision is acceptable in exchange for dramatic reductions in data size. Hence, these approaches play a crucial role in various domains, from signal processing and image compression to the emerging field of 3D data representation to compress complex datasets while maintaining a close approximation to the original. Especially, it is highly effective for compressing 3D Gaussian splatting (3DGS) data, where attributes like positions, colors, or spherical harmonics are often highly redundant and exhibit natural clustering tendencies.